OpenStudy (anonymous):

y=-2/3x+400 How do I graph this? Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

y=-2/3x+400 y= mx + b You start at b which is your Y Intercept of 400. Then you go down 2 and over 3 which is your slope.\